OpenStudy (anonymous):

two taps can fill an empty tank in 20 minutes and 30 minutes respectively. A third pipe can drain a full tank in 5 mins. If a tank is full, and all three pipes are opened together, how long will it take to empty the tank?

OpenStudy (anonymous):

20+30=50 minutes to fill it if it takes 5 minutes to drain it is 50/5 :D not so bad :)

OpenStudy (anonymous):

Hello @rebeccaskell94 sorry but that's the wrong answer.

OpenStudy (anonymous):

but thank you for your reply.

OpenStudy (anonymous):

D: I"m so sorry :( my bad

OpenStudy (apoorvk):

Let the tank's capacity be 'x'. So, in 1 minute, tap A fills x/20 of the tank, and tap B fills x/30th part of it. Now, both together fill in one minute, (x/20 + x/30)th of the tank = x/12 th of tank now, tap C (outlet) drains at the rate of x/5 per minute. So, when all 3 are open simultaneously, the net rate of flow when x/12 in and x/5 out, = x/5 - x/12 = 7x/60 per minute. So, if the net flow is '7x/60' 'out' per minute, then how much time to completely drain 'x' at this rate?

OpenStudy (anonymous):

60/7 minutes

OpenStudy (apoorvk):

righty-right!

OpenStudy (anonymous):

@apoorvk x/5 is negative work since it is draining the water out, so won't we write negative sign with x/5 and +ve sign with x/12?

OpenStudy (apoorvk):

It's the same thing. if 'outlet rate - inlet rate' is positive, that would mean the outlet has a larger rate, and there is a 'net' outward flow from the tank. If it's negative, it would mean the inlet rate is larger, and there is a net flow 'into' the tank. vice-versa, for cases when you write 'inlet - outlet'.